OpenStudy (anonymous):

Is the sequence 3, 12, 36, ... a geometric sequence? Explain

OpenStudy (saifoo.khan):

Use the formula to find out it's rate.

OpenStudy (anonymous):

what lol

OpenStudy (saifoo.khan):

\[\frac{12}{3} = 4\] \[\frac{36}{12}=3\] So their signs rate is different hence it means they aren't geometric.

OpenStudy (anonymous):

hey jenny dont mind me, this is for saifoo, saifoo please dont be mad at me:(.....................ok i made my point

OpenStudy (saifoo.khan):

The formula to find the rate is, \[\frac{\text{any term}}{\text{any term before it}}\]

OpenStudy (saifoo.khan):

@aminah.love18 , please check your inbox. :p

OpenStudy (anonymous):

need help...Is the sequence 5,9,15...an arithmetic sequence? explain